# Efficient Data Collection for Robotic Manipulation via Compositional Generalization

### v1

@@ Abstract
Data collection has become an increasingly important problem in robotic manipulation, yet there still lacks much understanding of how to effectively collect data to facilitate broad generalization. Recent works on large-scale robotic data collection typically vary a wide range of environmental factors during data collection, such as object types and table textures. While these works attempt to cover a diverse variety of scenarios, they do not explicitly account for the possible compositional abilities of policies trained on the data. If robot policies are able to compose different environmental factors of variation (e.g., object types, table heights) from their training data to succeed when encountering unseen factor combinations, then we can exploit this to avoid collecting data for situations that composition would address. To investigate this possibility, we conduct thorough empirical studies both in simulation and on a real robot that compare data collection strategies and assess whether visual imitation learning policies can compose environmental factors. We find that policies do exhibit composition, although leveraging prior robotic datasets is critical for this on a real robot. We use these insights to provide better practices for in-domain data collection by proposing data collection strategies that exploit composition, which can induce better generalization than naïve approaches for the same amount of effort during data collection. We further demonstrate that a real robot policy trained on data from such a strategy achieves a success rate of 77.5% when transferred to entirely new environments that encompass unseen combinations of environmental factors, whereas policies trained using data collected without accounting for environmental variation fail to transfer effectively, with a success rate of only 2.5%. We provide videos at our project website. [^1] http://iliad.stanford.edu/robot-data-comp/
@@ I Introduction
For robots to be practical and useful, they must robustly handle the wide variety of conditions they will encounter in the world. Recent advances in machine learning have shown that leveraging diverse, internet-scale data can be very effective in facilitating broad generalization [7, 38], as such data captures much of the complexity and variation in the domain that models will need to reason about. It therefore has been of great interest in the robotics community to apply a similar recipe to robotics, specifically using end-to-end imitation learning, which holds promise to effectively scale with large, diverse datasets to achieve broad generalization. However, we lack access to existing internet-scale robotics data, and furthermore, robotics data – especially human-collected demonstrations – is often limited and challenging to collect. While recent works have made significant progress in scaling real-world robotic manipulation datasets [44, 17, 35, 2], their scale is still much less than the datasets typically used for pre-training in computer vision and natural language processing. As a result, policies trained on such datasets still exhibit unsatisfactory zero-shot performance when deployed in new environments, often necessitating additional in-domain data collection for generalization. Given the lack of large-scale robotics data and the high cost of data collection, it is essential to focus on how to best collect this in-domain data.
@@
While prior work has studied generalization to unseen environmental factors in robotics [47, 46], there has not been as much focus on understanding compositional generalization in robotics, particularly for end-to-end imitation learning. Compositional generalization has been previously studied in other machine learning domains [19, 1, 25, 22, 23, 41, 4, 49], with results often suggesting that end-to-end neural models can struggle to achieve this. However, recent work has shown that large language models can exhibit strong compositional abilities, likely due to the diversity and scale of their training data [51, 13]. Ideally, prior datasets in robotics could also facilitate composition for robotics, but we still lack much understanding as to where these datasets benefit generalization.
Our key idea is that there is more guidance that can be provided for efficient data collection in robotics. While there has been a recent trend in scaling up robotic data collection, most prior works attempt to simply collect diverse data covering all possible combinations of different factors of variation [44, 17, 35, 2]. However, we hypothesize that imitation learning policies – similar to their large language model counterparts – can potentially exhibit compositional generalization, in particular when leveraging prior robotics datasets. In this work, we investigate this hypothesis, with the goal of prescribing better practices for robotic data collection.
@@ II Related Work
Generalization in Robot Learning. Much prior work has studied improving generalization in robot learning, such as by using pre-trained visual representations [36, 33, 30, 39, 11, 31, 20], or pre-training on diverse robot data [35, 34]. In our work, we focus on assessing compositional generalization of existing imitation learning methods. Some prior works have similarly studied the robustness of visual imitation learning models to different environmental factor shifts [47, 46, 8]. However, these works only consider generalization to out-of-distribution values for one or two factors at a time. Our work instead assesses compositional generalization, where policies must reason about unseen combinations of factor values.
@@
There has been some evidence of compositional generalization in end-to-end robotic imitation learning models. This includes composition of tasks in prior data with the conditions of a target domain [16], objects with manipulation skills [5, 43], semantic concepts from internet data with manipulation skills [6], and behaviors and semantic concepts across robot embodiments [5, 35]. However, our work provides more systematic and fine-grained analysis on when composition is possible, and considers a greater variety of environmental factors. Furthermore, unlike these works, we focus on using this analysis to guide better practices for data collection that take composition into account. We believe these differences allow our insights to be leveraged for more effective data collection in real-world settings.
@@ III-B Objective
We aim to train robotic manipulation policies using imitation learning that will robustly handle combinations of factor values from some desired distribution $p(f)$, by collecting demonstrations with associated factor values $f$ according to some strategy. We assume without loss of generality that $p(f)=\text{Uniform}(\mathcal{F}^{N})$, such that we desire to handle all possible combinations of factor values in $\mathcal{F}^{N}$. The most direct way to achieve this robustness is sufficiently covering all of $\mathcal{F}^{N}$, such as by sampling new values $f\sim\text{Uniform}(\mathcal{F}^{N})$ enough times, or even by collecting demonstrations for all $f\in\mathcal{F}^{N}$. However, as the cardinality of $\mathcal{F}^{N}$ increases exponentially with the number of factors $N$, collecting demonstrations for all possible factor value combinations is often impractical. Furthermore, constantly resampling new values $f\sim\text{Uniform}(\mathcal{F}^{N})$ can often be expensive or challenging in practice, as each factor change often requires some degree of manual effort.
Rather than attempting to collect data for all combinations of factor values in $\mathcal{F}^{N}$, it would be desirable to instead collect data with a focus on simply capturing all individual factor values, and then exploit compositional generalization of the learned policy to perform well on unseen combinations. With $N$ different factors and $k$ possible values each, collecting data covering all combinations would require $\mathcal{O}(k^{N})$ factor changes in the environment, while covering all individual factor values would only require $\mathcal{O}(kN)$, a dramatic reduction. However, this is a sensible option only if compositional generalization is actually achieved by the learned policy. As prior work has shown mixed results on the compositional abilities of end-to-end neural models, it remains important to investigate when compositional generalization happens in imitation learning for robotic manipulation. Therefore, to exploit compositional generalization for more efficient data collection, we must answer the following questions:
@@ III-C Data Collection Strategies
We describe the data collection strategies that we consider. We visualize these strategies for when the number of factors $N=2$ in Fig. 2, but these can easily be extended for $N>2$. The strategies involve periodically setting factor values in the environment during data collection. To quantify effort, we consider the total number of factor changes made in the environment. Note that changing multiple factors at a time involves multiple changes, e.g., setting the environment to have different values for all factors involves $N$ changes.
@@
Stair, L, and Diagonal are intended to exploit compositional generalization by prioritizing covering all individual factor values with the fewest amount of factor changes. We note that using factor changes as a measure of effort is a rough approximation. For example, this may not be accurate when it is more challenging to vary two factors together than separately. This could be the case if one had to collect data in multiple separate environments to vary a factor like table texture, but could collect data in a single environment to vary object type. In such situations, it may be easier to use the L strategy to collect data varying object type in only one environment, and collect data varying table texture in separate environments without varying object type, instead of varying both factors together like Stair and Diagonal would require.
@@ IV Simulation Experiments
Figure 4: Simulation results of data collection strategies for Pick Place. We compare strategies using the same amount of data for each. We evaluate every factor pair from 5 factors ($N=2$), average results across all pairs, and all 5 factors together ($N=5$). The strategies that exploit composition (Stair, L, Diagonal) generally outperform Random, and often approach the performance of Complete with enough factor changes. Stair generally performs the best, especially in the ($N=5$) setting. Error bars represent standard error across 5 seeds.
@@ IV-A Evaluation Protocol
We consider Pick Place and Door Open, two commonly studied tasks in robotics. We consider 5 factors, including 4 that generalization was shown to be challenging for in the original Factor World experiments (object texture, table texture, camera position, and distractor objects), as well as object position, as an example of a factor that more strongly affects required physical motions. We show example factor value combinations for table texture and object position for Pick Place in Fig. 3, and examples of individual factors for both tasks in Appendix A. To investigate which specific factors can be composed by policies effectively, we study the setting where $\mathcal{F}^{N}$ consists of one of the 10 possible factor pairs ($N=2$). To investigate composition in a more realistic setting where robustness to multiple factors is desired, we also study when $\mathcal{F}^{N}$ consists of all factors ($N=5$). For each factor, we sample $k=10$ values that the factor can take in $\mathcal{F}^{N}$. For each data collection strategy, we use a scripted expert policy to collect multiple datasets with different total amounts of factor changes, by setting new factor values according to the strategy at different rates. For example, with the Random strategy, 20 factor changes would involve resampling new factor combinations roughly twice as often as with 10 factor changes. We compare how the strategies scale with the total number of factor changes, which we use to quantify effort.
We only compare against Complete in the $N=2$ setting, as $N=5$ would require a minimum of $10^{5}$ demonstrations, which is impractical. Each dataset consists of 100 demonstrations for the $N=2$ experiments, and 1000 demonstrations for $N=5$. We train a policy on each dataset using behavior cloning, and evaluate on 100 episodes, each with a different $f$ randomly sampled from $\mathcal{F}^{N}$, to assess overall robustness to factor combinations. For the $N=2$ experiments, this consists of all possible $f\in\mathcal{F}^{N}$. We evaluate across 5 random seeds, which include different possible factor values in $\mathcal{F}^{N}$ for each. We provide more details on these experiments in Appendix A.
@@ IV-B Results
We illustrate our results for Pick Place in Fig. 4. We provide similar results for Door Open in Section A-A. We first focus on the $N=2$ setting. We see that No Variation (dashed gray) performs poorly overall, verifying that exposure to a variety of factor values is needed to generalize. The Single Factor strategies (shades of blue) also perform poorly for most factor pairs, verifying that most pairs require exposure to variation in both factors to generalize. Note that these strategies can vary only a single factor at most 10 times, leading to shorter lines in the plots than the others. We see that the strategies intended to exploit compositional generalization – Stair, L, Diagonal (shades of orange) – generally outperform Random (light purple) at all levels of factor changes. They also approach the performance of Complete (dashed green line) with only 20 factor changes (enough to capture all individual factor values), compared to the 100 changes needed for Complete. These results suggest that our policies exhibit strong pairwise compositional abilities for these factors, and that Stair, L, and Diagonal are able to exploit this for more efficient data collection. We also see that Stair does slightly better overall than L and Diagonal, possibly suggesting some benefit from a greater quantity and diversity of factor value combinations.
@@
In the $N=5$ setting, we see that performance is lower overall, indicating that generalization is harder when more factors are being varied. Stair, L, and Diagonal still outperform Random, but now Stair is the only strategy that is able to substantially improve over it. This suggests that a greater quantity and diversity of factor value combinations is more important for composition when there are more factors.
@@
Scaling with Dataset Size. In Fig. 5 (right), we investigate how generalization scales with dataset size. We focus on the $N=5$ setting, where generalization is more challenging, and therefore the impact of data scale is more pronounced. We report the success rate of policies trained on datasets from different strategies, which all have 50 factor changes (the amount needed for Stair, L, and Diagonal to capture all factor values), but different amounts of demonstrations. We also compare against No Variation, which does not scale well, verifying that simply more data is not enough to generalize. The other strategies do scale positively, but Stair does significantly better overall than the rest, and it is the only strategy that is able to significantly benefit when increasing from 500 to 1000 demonstrations. This suggests that having more data, even without greater factor diversity, can help facilitate improved compositional generalization, but having a greater quantity and diversity of factor combinations, which Stair provides, may be critical for enabling this scaling.
@@ V-A Evaluation Protocol
Robot Platform. We conduct experiments using a WidowX 250 6DOF robot arm. We adapt the hardware and control setup used in BridgeData V2 [44], except we directly mount the robot and over-the-shoulder RGB camera to a mobile table. While our hardware setup is not identical to the original, we still seek to leverage BridgeData V2 as prior data, which represents a realistic use case of a prior robotic dataset. To promote transfer, we tune our camera setup so that policies trained on only BridgeData V2 can sometimes perform basic pick-place tasks zero-shot, although we were unable to have it perform the exact tasks in our evaluation. We show our robot setup in Fig. 6, and provide more details in Section B-A.
@@
Task and Factors. We consider the task of placing a fork inside a container, on the countertop of an office kitchen. We primarily consider the following factors: object (fork) type, object (fork) position, container type, table height, and table texture. Each factor has $k=4$ possible values. We visualize our base values $f^{*}$, and also every deviation from $f^{*}$ by one factor, in Fig. 7. We additionally consider the factors camera position and distractor objects in secondary experiments.
@@
Experimental Setup. We adopt a different experimental setup for our real experiments, due to the much greater cost of collecting data, training policies, and evaluating. We collect datasets for only the $N=5$ setting with all of our primary factors. Because the objective of these experiments is to determine which factors can be composed by a real robot policy, we focus on data collection strategies intended to exploit composition. In particular, we evaluate Stair, because it performed the best in simulation, and L, because by collecting data for each factor separately, we can use this to study pairwise composition using a model trained on a single dataset. L may also be more practical in real-world situations where it is easier to collect data for factors separately. For these strategies, we collect 10 human demonstrations for each of the 16 combinations of factor values required to cover all values per factor, resulting in 160 total demonstrations for each strategy. We also compare against No Variation with the same number of demonstrations. We collect data exclusively in one kitchen, which we refer to as BaseKitch. We train policies either from scratch, or with the addition of BridgeData V2, to assess the impact of prior data on composition. We provide more details on our experimental setup in Section B-B.
Data Strategy L No Variation Train Method Bridge From Scratch Bridge Factor 1 Factor 2 Object Type Container Type Table Height Table Tex Object Type Container Type Table Height Table Tex Object Type Container Type Table Height Table Tex Object Position 8/9 5/9 2/9 5/9 0/9 0/9 0/9 1/9 3/9 2/9 1/9 1/9 Object Type 8/9 8/9 8/9 5/9 5/9 4/9 4/9 3/9 2/9 Container Type 5/9 6/9 4/9 6/9 2/9 3/9 Table Height 4/9 3/9 1/9 Overall 59/90 28/90 22/90
TABLE I: Real robot pairwise composition results. When leveraging BridgeData V2 as prior data, a policy is able to compose the factor values present in the L data to succeed on 59/90 compositional combinations of factor values. Without prior data, the model is unable to compose nearly as effectively, with compositional success rate dropping by roughly half. Prior data alone is also not enough to generalize to these situations, as a policy trained with prior data on No Variation also performs poorly.
Training. We adapt the diffusion goal-conditioned behavior cloning implementation from the BridgeData V2 code. For policies that use BridgeData V2 as prior data, unless otherwise stated, we first pre-train a policy on only the prior data, and then co-fine-tune on a mixture of in-domain and prior data. We verify that all policies succeed with base factor values $f^{*}$ in BaseKitch. We provide more details in Section B-C.

### v2

@@ Abstract
Data collection has become an increasingly important problem in robotic manipulation, yet there still lacks much understanding of how to effectively collect data to facilitate broad generalization. Recent works on large-scale robotic data collection typically vary many environmental factors of variation (e.g., object types, table textures) during data collection, to cover a diverse range of scenarios. However, they do not explicitly account for the possible compositional abilities of policies trained on the data. If robot policies can compose environmental factors from their data to succeed when encountering unseen factor combinations, we can exploit this to avoid collecting data for situations that composition would address. To investigate this possibility, we conduct thorough empirical studies both in simulation and on a real robot that compare data collection strategies and assess whether visual imitation learning policies can compose environmental factors. We find that policies do exhibit composition, although leveraging prior robotic datasets is critical for this on a real robot. We use these insights to propose better in-domain data collection strategies that exploit composition, which can induce better generalization than naïve approaches for the same amount of effort during data collection. We further demonstrate that a real robot policy trained on data from such a strategy achieves a success rate of 77.5% when transferred to entirely new environments that encompass unseen combinations of environmental factors, whereas policies trained using data collected without accounting for environmental variation fail to transfer effectively, with a success rate of only 2.5%. We provide videos at our project website. [^1] http://iliad.stanford.edu/robot-data-comp/
@@ I Introduction
For robots to be practical and useful, they must be robust to the wide variety of conditions they will encounter in the world. Recent advances in machine learning have shown that leveraging diverse, internet-scale data can be very effective in facilitating this kind of broad generalization [7, 40], as such data captures much of the complexity and variation in the domain that models will need to reason about. It therefore has been of great interest in the robotics community to apply a similar recipe to robotics, specifically using end-to-end imitation learning, which holds promise to effectively scale with large, diverse datasets to achieve broad generalization. However, we lack access to existing internet-scale robotics data, and furthermore, robotics data – especially human-collected demonstrations – is often limited and challenging to collect. While recent works have made significant progress in scaling real-world robotic manipulation datasets [47, 17, 36, 2], their scale is still much less than the datasets typically used for pre-training in computer vision and natural language processing. As a result, policies trained on such datasets often still exhibit unsatisfactory zero-shot performance when deployed in new environments, necessitating additional in-domain data collection for generalization. Given the lack of large-scale robotics data and the high cost of data collection, it is essential to focus on how to best collect this in-domain data.
@@
While prior work has studied generalization to unseen environmental factors in robotics [50, 49, 39], there has not been as much focus on understanding compositional generalization in robotics, particularly for end-to-end imitation learning. Compositional generalization has been previously studied in other machine learning domains [20, 1, 26, 23, 24, 44, 4, 52], with results often suggesting that end-to-end neural models can struggle to achieve this. However, recent work has shown that large language models can exhibit strong compositional abilities, likely due to the diversity and scale of their training data [54, 13]. Ideally, prior datasets in robotics could also facilitate composition for robotics, but we currently lack much understanding as to where these datasets benefit generalization.
Our key idea is that understanding composition can provide more guidance for efficient data collection in robotics. While there has been a recent trend in scaling up robotic data collection, most prior works attempt to simply collect diverse data covering all possible combinations of different factors of variation, without explicitly accounting for composition [47, 17, 36, 2]. However, we hypothesize that imitation learning policies – similar to their large language model counterparts – can potentially exhibit compositional generalization, in particular when leveraging prior robotics datasets. In this work, we investigate this hypothesis, with the goal of prescribing better practices for robotic data collection.
@@ II Related Work
Generalization in Robot Learning. Much prior work has studied improving generalization in robot learning, such as by using pre-trained visual representations [37, 34, 31, 41, 11, 32, 21], or leveraging diverse robot data [36, 35]. In our work, we focus on assessing compositional generalization of existing imitation learning methods. Some prior works have similarly studied the robustness of visual imitation learning to environmental factor shifts [50, 49, 8, 39]. However, these works consider generalization to unseen factor values, and primarily focus on how factors individually affect generalization, without extensive focus on how factors interact. Our work instead assesses compositional generalization, where policies must reason about unseen combinations of factor values.
@@
There has been some evidence of compositional generalization in end-to-end robotic imitation learning. This includes composition of tasks in prior data with the conditions of a target domain [16], objects with manipulation skills [5, 46], semantic concepts from internet data with manipulation skills [6], and behaviors and semantic concepts across robot embodiments [5, 36]. However, our work provides more systematic and fine-grained analysis on when composition is possible, and considers a greater variety of environmental factors. Furthermore, unlike these works, we focus on using this analysis to guide better practices for data collection that take composition into account. We believe these key differences allow our insights to be leveraged for more effective data collection in real-world settings.
@@ III-B Objective
We aim to train robotic manipulation policies using imitation learning that will robustly handle combinations of factor values from some desired distribution $p(f)$, by collecting demonstrations with associated factor values $f$ according to some strategy. We assume without loss of generality that $p(f)=\text{Uniform}(\mathcal{F}^{N})$, such that we desire to handle all factor value combinations in $\mathcal{F}^{N}$. The most direct way to achieve this is sufficiently covering all of $\mathcal{F}^{N}$, such as by sampling new values $f\sim\text{Uniform}(\mathcal{F}^{N})$ enough times, or even by collecting demonstrations for all $f\in\mathcal{F}^{N}$. However, as the cardinality of $\mathcal{F}^{N}$ increases exponentially with the number of factors $N$, collecting demonstrations for all possible factor value combinations is often impractical. Furthermore, constantly resampling new values $f\sim\text{Uniform}(\mathcal{F}^{N})$ can often be expensive or challenging in practice, as each factor change often requires some degree of manual effort.
Rather than attempting to collect data for all factor value combinations in $\mathcal{F}^{N}$, it would be desirable to instead collect data with a focus on capturing individual factor values, and then exploit compositional generalization of the learned policy to perform well on unseen combinations. With $N$ factors and $k$ possible values each, collecting data for all combinations would require $\mathcal{O}(k^{N})$ factor changes in the environment, while covering all individual factor values would only require $\mathcal{O}(kN)$ changes, a dramatic reduction. However, this is a sensible option only if composition is actually achieved by the learned policy. As prior work has shown mixed results on the compositional abilities of end-to-end neural models, it remains important to investigate when compositional generalization happens in imitation learning for robotic manipulation. Therefore, to exploit compositional generalization for more efficient data collection, we must answer the following questions:
@@ III-C Data Collection Strategies
We describe the data collection strategies that we consider. We visualize them for when the number of factors $N=2$ in Fig. 2, but these can easily be extended for $N>2$. We also provide pseudocode for Diagonal, L, and Stair in Section A-A. The strategies involve periodically setting factor values in the environment during data collection. To quantify effort, we consider the total number of factor changes made in the environment. Note that changing multiple factors at a time involves multiple changes, e.g., setting the environment to have different values for all factors involves $N$ changes.
@@
Stair, L, and Diagonal are intended to exploit compositional generalization by prioritizing covering individual factor values with the fewest amount of factor changes. We note that using factor changes as a measure of effort is a rough approximation. For example, this may not be accurate when it is more challenging to vary two factors together than separately. This could be the case if one had to collect data in multiple separate environments to vary a factor like table texture, but could collect data in a single environment to vary object type. In such situations, it may be easier to use the L strategy to collect data varying object type in only one environment, and collect data varying table texture in separate environments without varying object type, instead of varying both factors together like Stair and Diagonal would require.
@@ IV Simulation Experiments
Figure 4: Simulation results of data collection strategies for Pick Place. We report results where $\mathcal{F}^{N}$ consists of each possible factor pair ($N=2$), average results across all pairs, and results where $\mathcal{F}^{N}$ consists of all factors ($N=5$). All points within the same subplot use the same amount of demonstrations. The strategies that exploit composition (Stair, L, Diagonal) generally outperform Random, and often approach Complete. Stair generally performs the best, especially in the $N=5$ setting. Error bars represent standard error across 5 seeds.
@@ IV-A Evaluation Protocol
We consider Pick Place and Door Open, two commonly studied tasks in robotics. We consider 5 factors, including 4 that generalization was challenging for in the original Factor World experiments (object texture, table texture, camera position, and distractor objects), as well as object position, as an example of a factor that more strongly affects required physical motion. We show example combinations of table texture and object position for Pick Place in Fig. 3, and examples of individual factor values for both tasks in Appendix B. To study which specific factors can be composed by policies effectively, we consider the settings where $\mathcal{F}^{N}$ consists of one of the 10 possible factor pairs ($N=2$). To study composition in a more realistic setting where robustness to more factors is desired, we also consider when $\mathcal{F}^{N}$ consists of all factors ($N=5$). For each factor, we sample $k=10$ values that the factor can take in $\mathcal{F}^{N}$. For each data strategy, we use a scripted expert policy to collect multiple datasets with different total amounts of factor changes, by setting new factor values according to the strategy at different rates. For example, 20 factor changes would involve setting new values roughly twice as often as with 10 changes. When setting new values, we choose which values to set at random from what is permissible by the current strategy. We compare how the strategies scale with the total number of factor changes, which we use to quantify effort.
We only compare against Complete in the $N=2$ setting, as $N=5$ would require a minimum of $10^{5}$ demonstrations, which is impractical. Each dataset consists of 100 demonstrations for the $N=2$ experiments, and 1000 demonstrations for $N=5$. We train a policy on each dataset using behavior cloning, and evaluate on 100 episodes, each with a different $f$ randomly sampled from $\mathcal{F}^{N}$, to assess overall robustness to factor combinations. For the $N=2$ experiments, this consists of all possible $f\in\mathcal{F}^{N}$. We evaluate across 5 random seeds, which include different possible factor values in $\mathcal{F}^{N}$ for each seed. We provide more details in Appendix B.
@@ IV-B Results
We illustrate our results for Pick Place in Fig. 4. We first focus on the $N=2$ setting. We see that No Variation (dashed gray) performs poorly, verifying that a variety of factor values is needed to generalize. The Single Factor strategies (shades of blue) also perform poorly for most factor pairs, verifying that most pairs require variation in both factors to generalize. Note that these strategies can vary only a single factor at most 10 times, leading to shorter lines in the plots. We see that the strategies intended to exploit composition – Stair, L, Diagonal (shades of orange) – generally outperform Random (light purple) at all levels of factor changes. They also approach the performance of Complete (dashed green line) with only 20 factor changes (enough to capture all individual factor values), compared to the 100 changes needed for Complete. These results suggest that these policies exhibit strong pairwise compositional abilities for these factors, and that Stair, L, and Diagonal are able to exploit this for more efficient data collection. We also see that Stair does slightly better overall than L and Diagonal, possibly suggesting some benefit from a greater quantity and diversity of factor value combinations.
@@
In the $N=5$ setting, performance is lower overall, indicating that generalization is harder when more factors are varied. Furthermore, Stair is now the only strategy that substantially improves over Random. This suggests that a greater quantity and diversity of factor value combinations is more important for composition when there are more factors involved.
We provide similar results for Door Open, results with different data augmentation, and results on accounting for factor value similarity during data collection in Appendix B.
@@
Scaling with Dataset Size. In Fig. 5 (right), we investigate how generalization scales with dataset size. We focus on the $N=5$ setting, where generalization is more challenging, and therefore the impact of data scale is more pronounced. We report the success rate of policies trained on datasets from different strategies, which all have 50 factor changes (the amount needed for Stair, L, and Diagonal to capture all factor values), but different amounts of demonstrations. We also compare against No Variation, which does not scale well, verifying that simply more data is not enough to generalize. The other strategies do scale positively, but Stair does significantly better than the rest, and it is the only strategy that is able to significantly benefit when increasing from 500 to 1000 demonstrations. This suggests that having more data, even without greater factor diversity, can help facilitate improved compositional generalization, but having a greater quantity and diversity of factor combinations, which Stair provides, may be critical for enabling this scaling.
@@ V-A Evaluation Protocol
Robot Platform. We conduct experiments using a WidowX 250 6DOF robot arm. We adapt the hardware and control setup used in BridgeData V2 [47], except we directly mount the robot and over-the-shoulder RGB camera to a mobile table. While our hardware setup is not identical to the original, we still seek to leverage BridgeData V2 as prior data, which represents a realistic use case of a prior robotic dataset. To promote transfer, we tune our camera setup so that policies trained on only BridgeData V2 can sometimes perform basic pick-place tasks zero-shot, although we were unable to have it perform the exact tasks in our evaluation. We show our robot setup in Fig. 6, and provide more details in Section C-A.
@@
Task and Factors. We consider the task of placing a fork inside a container, on the countertop of an office kitchen. We primarily consider the following factors: object (fork) type, object (fork) position, container type, table height, and table texture. Each factor has $k=4$ possible values. We visualize our base values $f^{*}$, and also every deviation from $f^{*}$ by one factor, in Fig. 7. We additionally consider the factors object/container position, object (fork) orientation, distractor objects, and camera position in secondary experiments.
@@
Experimental Setup. We adopt a different experimental setup for our real experiments, due to the much greater cost of collecting data, training policies, and evaluating. We collect datasets for only the $N=5$ setting with all of our primary factors. Because the objective of these experiments is to determine which factors can be composed by a real robot policy, we focus on data collection strategies intended to exploit composition. In particular, we evaluate Stair, because it performed the best in simulation, and L, because by collecting data for each factor separately, we can use this to study pairwise composition using a model trained on a single dataset. L may also be more practical in real-world situations where it is easier to collect data for factors separately. For these strategies, we collect 10 human demonstrations for each of the 16 combinations of factor values required to cover all values per factor, resulting in 160 total demonstrations for each strategy. We also compare against No Variation with the same number of demonstrations. We collect data exclusively in one kitchen, which we refer to as BaseKitch. We train policies either from scratch, or with the addition of BridgeData V2, to assess the impact of prior data on composition. We provide more details on our experimental setup in Section C-B.
Data Strategy L No Variation Train Method Bridge From Scratch Bridge Factor 1 Factor 2 Object Type Container Type Table Height Table Tex Object Type Container Type Table Height Table Tex Object Type Container Type Table Height Table Tex Object Pos 8/9 5/9 2/9 5/9 0/9 0/9 0/9 1/9 3/9 2/9 1/9 1/9 Object Type 8/9 8/9 8/9 5/9 5/9 4/9 4/9 3/9 2/9 Container Type 5/9 6/9 4/9 6/9 2/9 3/9 Table Height 4/9 3/9 1/9 Overall 59/90 28/90 22/90
TABLE I: Real robot pairwise composition results for our “put fork in container” task. When leveraging BridgeData V2 as prior data, a policy is able to compose the factor values present in the L data to succeed on 59/90 compositional combinations of factor values. Without prior data, the model is unable to compose nearly as effectively, with compositional success rate dropping by roughly half. Prior data alone is also not enough to generalize to these situations, as a policy trained with prior data on No Variation also performs poorly.
Training. We primarily train policies based on the diffusion goal-conditioned behavior cloning method proposed in BridgeData V2 [47], adapting the implementation provided in their code. For policies that use BridgeData V2 as prior data, unless otherwise stated, we first pre-train a policy on only the prior data, and then co-fine-tune on a mixture of in-domain and prior data. We verify that all policies succeed with base factor values $f^{*}$ in BaseKitch. We provide more details in Section C-C.